OpenStudy (anonymous):

WHAT IS (3/16)(2/5)DIVDIDE (7/4)(3/2)

OpenStudy (anonymous):

1/35

OpenStudy (anonymous):

Ya it's 1/35

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[(3/16) (2/5)/ (7/4) (3/2) \]\[\left.\frac{3}{40}\right/ (7/4)(3/2) \]\[\left(\frac{3}{70}\right)(3/2) \]\[\frac{9}{140} \]